# Grasp as You Say: Language-guided Dexterous Grasp Generation

### v1

@@ 5.3 Comparison with SOTA methods
The comparison results are presented in Table 1. We reproduce the SOTA methods to suit our task by concatenating the language condition with the point cloud features. As seen in the Table, our DexGYSGrasp framework significantly outperforms all previous methods in terms of intention consistency and grasp diversity, while also achieving comparable performance in grasp quality. Previous methods struggle with learning a robust language conditional grasp distribution due to the optimization challenges outlined in Section 4.1. They often yield misaligned yet high quality grasps, resulting in comparable grasp quality, but less aligned intention and limited diversity compared to our framework. Overall, these results confirm that our framework achieves SOTA performance in generating intention-aligned, high-quality and diverse grasps.
@@
5.4 Necessity of Progressive Components, and Losses
@@
setp1 setp2 setp3 $Con.\downarrow$ $Q_{1}\uparrow$ $Pen\downarrow$
@@ 5.5 Plug-and-play Experiments
We conducted experiments to evaluate the applicability of our insights to other state-of-the-art (SOTA) methods. Specifically, we trained GraspCAVE and SceneDiffuser without the object penetration constraint and trained the quality grasp component (QGC) to refine the coarse outcomes. As depicted in Figure 4, removing the object penetration loss leads to improved intention consistency, which corroborates our findings discussed in Section 4.1. Moreover, our quality grasp component can significantly enhance grasp quality while maintaining the intention consistency.
@@ 5.6 Effectiveness of Hand-Object Interaction Retargeting
We conducted ablation studies to evaluate our Hand-Object Interaction Retargeting (HOIR) strategy in constructing DexDYSNet dataset. As shown in Table 4, our three-step HOIR significantly improves both the quality and the intention consistency progressively. We observed that optimizing all losses in Equations 1 and 2 in one step (all in one stage), results in worse contact consistency and better grasp quality. Similar outcomes occur when the root translation is not fixed in step 3 (w/o fix translation). We believe this trade-off arises from inherent noise in the hand-object interaction data and the structural differences between human grasps and dexterous hands, making it challenging to excel in all aspects. Overall, we think that three-step HOIR strategy achieves more comprehensive outcomes, especially in the most important aspect of hand object contact consistency.
@@ 5.7 Experiments in Real World
We conducted real-world grasp experiments to verify the practical application of our methods, as shown in Figure 7. The experiments are conducted on an Allegro hand, a Flexiv Rizon 4 arm and an Intel Realsense D415 camera. Although our framework is designed for full object point clouds, we integrate several off-the-shelf methods to enhance its practicality. Specifically, partial object point clouds are obtained through visual grounding liu2023groundingdino and SAM kirillov2023SAM , which are then fed into a point cloud completion network yuan2018pcn to obtain full point clouds. In execution, we first move the arm to the 6-DOF pose of the dexterous hand root node, and then control the dexterous hand joint angles to the predicted poses. Real world experiments further validate the effectiveness of our method.
@@ 6 Conclusions
We believe that enabling robots to perform high quality dexterous grasps aligned with human language is crucial within the deep learning and robotics communities. In this paper, we explore this novel task, term "Dexterous Grasp as You Say". This task is non-trival, we propose a DexGYSNet dataset and a DexGYSGrasp framework to accomplish it. DexGYSNet dataset is constructed cost-effectively using the object-hand interaction retargeting strategy and the language guidance annotation system assisted by LLMs. Building on DexGYSNet, DexGYSGrasp framework, comprised of two progressive components, which can achieve intention-aligned, high diversity, and high quality dexterous grasp generation. Extensive experiments in DexDYSNet and real-world settings demonstrate that our framework significantly outperforms all SOTA methods, confirming the potential and effectiveness of our approach.

### v2

@@ 5.3 Comparison with SOTA methods
The comparison results are presented in Table 1. We reproduce the SOTA methods to suit our task by concatenating the language condition with the point cloud features, the details can be found in Appendix A.3.3. As seen in the Table, our framework significantly outperforms all previous methods in terms of intention consistency and grasp diversity, while also achieving comparable performance in grasp quality. Previous methods struggle with learning a robust language conditional grasp distribution due to the optimization challenges outlined in Section 4.1. They often yield misaligned yet high quality grasps, resulting in comparable grasp quality, but less aligned intention and limited diversity compared to our framework. Overall, these results confirm that our framework achieves SOTA performance in generating intention-aligned, high-quality and diverse grasps.
@@
5.4 Necessity of Progressive Components and Losses
@@
step1 step2 step3 $Con.\downarrow$ $Q_{1}\uparrow$ $Pen\downarrow$
@@ 5.5 Plug-and-play Experiments
We conducted experiments to evaluate the applicability of our insights to other state-of-the-art (SOTA) methods. Specifically, we trained GraspCAVE and SceneDiffuser without the object penetration constraint and trained the quality grasp component (QGC) to refine the coarse outcomes. As depicted in Table 4, removing the object penetration loss leads to improved intention consistency, which corroborates our findings discussed in Section 4.1. Moreover, our quality grasp component can significantly enhance grasp quality while maintaining the intention consistency.
@@ 5.6 Effectiveness of Hand-Object Interaction Retargeting
We conducted ablation studies to evaluate our Hand-Object Interaction Retargeting (HOIR) strategy in constructing DexGYSNet dataset. As shown in Table 4, our three-step HOIR significantly improves both the quality and the intention consistency progressively. We observed that optimizing all losses in Equations 1 and 2 in one step (all in one stage), results in worse contact consistency and better grasp quality. Similar outcomes occur when the root translation is not fixed in step 3 (w/o fix translation). We believe this trade-off arises from inherent noise in the hand-object interaction data and the structural differences between human grasps and dexterous hands, making it challenging to excel in all aspects. Overall, we think that three-step HOIR strategy achieves more comprehensive outcomes, especially in the most important aspect of hand object contact consistency.
Figure 7: Visualization of real world experiments.
@@ 5.7 Experiments in Real World
We conducted real-world grasp experiments to verify the practical application of our methods, as shown in Figure 7. The experiments are conducted on an Allegro hand, a Flexiv Rizon 4 arm and an Intel Realsense D415 camera. Although our framework is designed for full object point clouds, we integrate several off-the-shelf methods to enhance its practicality. Specifically, partial object point clouds are obtained through visual grounding liu2023groundingdino and SAM kirillov2023SAM , which are then fed into a point cloud completion network yuan2018pcn to obtain full point clouds. In execution, we first move the arm to the 6-DOF pose of the dexterous hand root node, and then control the dexterous hand joint angles to the predicted poses. Real world experiments further validate the effectiveness of our method. More implementation details can be found in Appendix A.5.
@@ 6 Conclusions
We believe that enabling robots to perform high quality dexterous grasps aligned with human language is crucial within the deep learning and robotics communities. In this paper, we explore this novel task, “Dexterous Grasp as You Say” (DexGYS). This task is non-trival, we propose a DexGYSNet dataset and a DexGYSGrasp framework to accomplish it. DexGYSNet dataset is constructed cost-effectively using the object-hand interaction retargeting strategy and the language guidance annotation system assisted by LLMs. Building on DexGYSNet, DexGYSGrasp framework, comprised of two progressive components, which can achieve intention-aligned, high diversity, and high quality dexterous grasp generation. Extensive experiments in DexGYSNet and real-world settings demonstrate that our framework significantly outperforms all SOTA methods, confirming the potential and effectiveness of our approach.
Acknowledgements
This work was supported partially by NSFC(92470202, U21A20471), Guangdong NSF Project (No. 2023B1515040025). Additionally, I sincerely thank the help of Guo-Hao Xu and Dian Zheng for the valuable suggestions for the paper.
Appendix A Appendix / supplemental material
A.1 DexGYSGrasp Details
A.1.1 Diffusion Background
The diffusion model is used in our intention and diversity grasp component to generate grasp distribution with aligned intention and high diversity, which represents a class of generative models characterized by a forward process of noise addition and a reverse process of denoising. The forward process entails a Markov Chain that incrementally introduces Gaussian noise into the data across multiple time steps. Originating from the initial data $x_{0}$, this process transitions the data to conform with a standard Gaussian distribution $x_{T}$ after $T$ time steps. This transformation is mathematically formulated as follows:
$x_{t}=\sqrt{\alpha_{t}}x_{t-1}+\sqrt{1-\alpha_{t}}\epsilon_{t-1},$ (6)
where $\alpha_{t}$ denotes a time-dependent noise coefficient, $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{i}$.
Therefore, $x_{t}$ given $x_{0}$ follows a normal distribution,
$q(x_{t}|x_{0})=\mathcal{N}(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\mathbf{I}).$ (7)
The first equation delineates the stepwise diffusion, whereas the second equation offers a direct approximation of any intermediate state $x_{t}$ from $x_{0}$.
During the reverse process, the model is trained to closely approximate the reverse conditional distribution $p(x_{t-1}|x_{t})$, which is described as:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\sigma^{2}_{\theta}(x_{t},t)),$ (8)
where $\mu_{\theta}(x_{t},t)$ and $\sigma^{2}_{\theta}(x_{t},t)$ are the mean and variance parameters for the distribution of $x_{t-1}$, respectively, and $\theta$ indicates the parameters of the model used to predict $\epsilon$ from $x_{t}$.
The classical sampling strategy for the reverse process is exemplified by DDPM ho2020DDPM , where the model iteratively learns to reverse the noise addition process to reconstruct the original data from noise. It estimates the distribution $p(x_{t-1}|x_{t})$ and predicts the noise $\epsilon$, represented by:
$\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)$ $=\tilde{\boldsymbol{\mu}}_{t}\left(\mathbf{x}_{t},\frac{1}{\sqrt{\bar{\alpha}_{t}}}\left(\mathbf{x}_{t}-\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t}\right)\right)\right)$ (9)
$=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}\left(\mathbf{x}_{t},t\right)\right),$ (10)
$x_{t-1}$ $=\boldsymbol{\mu}_{\theta}\left(\mathbf{x}_{t},t\right)+\sigma_{\theta}(x_{t},t)z,$ (11)
where $\sigma_{\theta}$ consists of non-trainable, time-dependent constants, and $z$ represents Gaussian noise.
A.1.2 Intention and Diversity Grasp Component
Point Encoder We utilize a three-layer PointNet++ pointnet++ as our point encoder, following recent works graspness ; dexgraspnet ; gendexgrasp ; xu2024dgtr in the field of robotic grasping. Specifically, each layer $l_{i},i\in{1,2,3}$, receives point clouds and corresponding features (initially the raw XYZ coordinates for the first layer) from the preceding layer. It then performs down-sampling and feature aggregation using the "set-aggregation" operation pointnet++ . The aggregated features are processed by a three-layer perceptron, which consists of three $Linear-BatchNorm-ReLU$ blocks. The output of point encoder is $\mathcal{F}_{obj}\in\mathbb{R}^{N_{obj}\times C_{obj}}$.
Language Encoder For the language encoder, we employ the CLIP model with the ViT-L/14 architecture radford2021clip . The input text sequence is tokenized and converted into token embeddings with positional embeddings added. This sequence is processed through multiple Transformer encoder layers to obtain the language feature $\mathcal{F}_{lan}\in\mathbb{R}^{N_{lan}\times C_{lan}}$.
Transformer Decoder We employ four layers of MLPs and Transformer as decodertransformer ; scene_diffuser . The time embedding and pose feature are incorporated in MLPs to obtain $\mathcal{F}_{dex\_t}$. Subsequently, the $\mathcal{F}_{dex\_t}$ serves as the query, and the concatenated features of language and object, $\mathcal{F}_{lan\_obj}$, act as key and value in Transformer block. The corss attention process is formalized as:
$\mathcal{F}_{out}=\text{softmax}\left(\frac{f_{q}(\mathcal{F}_{dex\_t})f_{k}(\mathcal{F}_{lan\_obj})^{T}}{\sqrt{d_{k}}}\right)f_{v}(\mathcal{F}_{lan\_obj}),$ (12)
where ${f}_{q},{f}_{k},{f}_{v}$ are MLPs, and $d_{k}$ is the channel of features. Finally, a MLP is adopted to regress the dexterous grasp parameters $\hat{\mathcal{G}}^{dex}$. $\mathcal{F}_{dex\_t}={f}_{0}(f_{1}({F}_{dex})+f_{2}(\mathcal{F}_{time}))$, where ${f}_{0},{f}_{1},{f}_{2}$ are MLPs.
A.1.3 Quality Grasp Component
Quality grasp component tasks coarse pose $\hat{\mathcal{G}}^{dex}$, coarse hand point clouds $\mathcal{H}(\hat{\mathcal{G}}^{dex})$ and object point clouds $\mathcal{O}$ as input, and outputs refined grasp $\tilde{\mathcal{G}}^{dex}$. The object and hand are encoded by the PointNet++ same with intention and diversity grasp component. And in the transformer decoder, coarse pose features act as the query, object and hand features serve as key and value.
Figure 8: Inference pipeline of our DexGYSGrasp.
A.1.4 Inference Pipeline
We also demonstrate the inference pipeline of our DexGYSGrasp, as shown in Fig. 8. We sample a random noise from Gaussian distribution as the input, with the point cloud and language guidance as the conditions. We first generate the coarse grasp by the intention and diversity grasp component, and then refine it with the quality grasp component.
A.1.5 Loss Function
This section provides a detailed exposition of the loss functions utilized during the construction of datasets and the training of models.
Parameter Regression Loss. We utilize the mean squared error (MSE) to quantify the deviation between the generated dexterous hand pose $\hat{\mathcal{G}}^{dex}$ and the ground truth $\mathcal{G}^{dex}$.
$\mathcal{L}_{para}=\frac{1}{N}\sum_{i=1}^{N}\|\mathcal{G}^{dex,i}-\hat{\mathcal{G}^{dex,i}}\|_{2}^{2}.$ (13)
Hand Chamfer Loss. The predicted hand point clouds $\mathcal{H}(\hat{\mathcal{G}}^{dex})$ and the ground truth $\mathcal{H}(\mathcal{G}^{dex})$ are derived by sampling from the hand mesh. We then compute the chamfer distance to assess the discrepancies between the predicted and ground-truth hand shapes.
$\mathcal{L}_{chamfer}=\sum_{x\in\mathcal{H}(\mathcal{G}^{dex})}\min_{y\in\mathcal{H}(\hat{\mathcal{G}}^{dex})}\|x-y\|_{2}^{2}+\sum_{x\in\mathcal{H}(\hat{\mathcal{G}}^{dex})}\min_{y\in\mathcal{H}(\mathcal{G}^{dex})}\|x-y\|_{2}^{2}.$ (14)
Contact Map Loss. The contact map loss $\mathcal{L}_{cmap}$ ensures consistency between the predicted hand contact map $\hat{c}^{obj}$ on object and the target $c^{obj}$. The contact map is calculate by the distance from object point to the closest dexterous hand point.
$\mathcal{L}_{cmap}=\sum_{i}\|c_{i}^{obj}-\hat{c}_{i}^{obj}\|_{2}^{2}.$ (15)
Object Penetration Loss. The object penetration loss $\mathcal{L}_{pen}$ penalizes the depth of hand-object penetration, where $d_{i}^{sdf}$ denotes the signed distance from the object point to the hand mesh.
$\mathcal{L}_{pen}=\sum_{i}\mathbb{I}(d_{i}^{sdf}>0)\cdot d_{i}^{sdf}.$ (16)
Self-Penetration Loss. The self-penetration loss $\mathcal{L}_{spen}$ punishes the penetration among the different parts of the hand, where $p^{dex,sp}$ denotes predefined anchor spheres on the hand dexgraspnet .
$\mathcal{L}_{spen}=\sum_{i,j}\mathbb{I}(i\neq j)\cdot\max(\delta-d(p_{i}^{dex,sp},p_{j}^{dex,sp})).$ (17)
Joint Angle Loss. Given the physical structure limitations of the robotic hand, each joint has designated upper and lower limits. The joint angle loss penalizes deviations from these limits.
$\mathcal{L}_{joint}=\sum_{i}(\max(q_{i}-q_{i}^{max})+\max(q_{i}^{min}-q_{i})).$ (18)
Figure 9: The Extension of DexGYSNet to more dexterous hands.
A.2 DexGYSNet Datasets Details
A.2.1 Prompt of LLM
We introduce the prompt for using GPT-3.5 in this section.
System Prompt: "You are an assistant in creating language instruction, aimed at guiding robot on how to grasp objects. Given a brief instruction and a fine-gained interaction information. Your task is generate a natural and more informative instruction. The instruction should start with the given brief instruction, which is limited in a sentence and about 10-15 words."
User Prompt: "Brief instruction: To <brief intention> a <object category>. Hand-object interaction information: <contact information>. "
The brief intention and object category are sourced from the hand-object dataset OakInk yang2022oakink . The contact information is derived by calculating the distances from predefined contact anchors on each finger to the segmentation parts of the object. Details on predefined contact anchors are available in DexGraspNet dexgraspnet , and segmentations are annotated in OakInk yang2022oakink .
An example of user prompt is: "Brief instruction: To use a trigger sprayer. Hand-object interaction information: forefinger touches the trigger. thumb, middle finger, ring finger and little finger touches the finger." An example of LLM output is: "To use a trigger sprayer, press the trigger with your forefinger and hold the bottle with your other fingers."
A.2.2 DexGYSNet Extension
Our cost-effective dataset construction strategy can be easily extended to various types of dexterous hands. As shown in Figure 9, besides the Shadow Handshadowhand , which features a highly biomimetic design replicating most degrees of freedom of human hands at a high cost of \$100,000, we also expand to the Allegro Hand allegrohand and Leap Hand shaw2023leap . These latter models, while offering fewer degrees of freedom, are significantly more affordable, costing \$16,000 and \$2,000, respectively, making them practical for promoting the use of robotic arms in laboratory environments. We have trained our method on the DexGYSNet dataset using the Allegro Hand and implemented it in real robot experiments.
Figure 10: (a) Evaluation of intention consistency using the Fréchet Inception Distance between the <generation hand and object> and the ground truth. (b) When the ground truth is not available (e.g., evaluation on a 3D object dataset), we employ GPT4-o for evaluation.
A.3 Implementation Details
A.3.1 Dataset Split
We split the DexDYS dataset at the level of object instances. Specially, for all objects within each category, 80% of the objects instances are used for training and 20% for evaluation. Concretely, the training set includes approximately 1,200 objects with 40k grasps, while the evaluation set comprises about 300 objects with 10k grasps. Therefore, all objects in the test set of DexGYSNet don’t exist in the training set.
A.3.2 Metrics Detials
Target Assignment. For target assignment in the testing phase, the grasp targets of an object-guidance pair consist of all poses that share the same contact part and brief guidance. And the matrices of intention consistency are calculated by comparing the prediction to the most similar grasp target.
Fréchet Inception Distance, which is commonly used in generative taskguo20223Dhuman by measuring the distance between the generated distribution and the ground truth distribution. We use sampling point cloud features extracted fromnichol2022pointe to calculate P-FID and rendering image features extracted fromheusel2017fid to calculate FID. The details are shown in Figure 10 (a).
Chamfer distance, denoted as $CD$, is used to measure the distance between predicted hand point clouds and targets to measure the consistency from the aspect of hand consistency. Please look at Equation 14 for details.
Contact distance, denoted as $Con.$ to measure the L2 distance of object contact map between the prediction and targets to measure the consistency from the aspect of object contact consistency. Please look at Equation 15 for details.
Success rate. We evaluate the grasp success rate in Issac Gym simulation environment. To simulate the force exerted by dexterous hands grasping objects in real environments, we contract each finger in the direction of the object. If the grasp can withstand at least one of the six directions of gravity, it is considered successful.
Mean $Q_{1}$. Intuitively, the $Q_{1}$ metric reflects the norm of the smallest wrench which can disrupt the stability of a grasp. We follow dexgraspnet to set the contact threshold to 1cm and set the penetration threshold to 5mm. Any grasp with its maximal penetration depth greater than 5mm is considered invalid and we set the $Q_{1}$ of it to 0 before taking the average.
Maximal penetration depth, which is the maximal penetration depth from the object point cloud to hand meshes.
Diversity. We use the standard deviation of translation $\delta_{t}$, rotation $\delta_{r}$, and joint angle $\delta_{q}$ to measure the diversity of generated grasps. We perform eight samples in the intention and diversity component under the same input conditions, and each sample is individually sent to the quality component for refinement. Before calculation, $\delta_{r}$ and $\delta_{q}$ are converted to Euler angles in degrees, while $\delta_{t}$ is measured in centimeters.
A.3.3 Implementation Details of SOTA Methods
We replicate SOTA methods on our DexGYSNet dataset using the same encoder structure and the loss functions defined in Equation 5 to ensure fair comparison. Specifically, we reimplement GraspCAVE based on cvae , GraspTTA grasptta , SceneDiffuser scene_diffuser , and DGTR xu2024dgtr . To introduce language information, we use an identical CLIP language encoder. For GraspCAVE, we concatenate the language feature, object feature, and latent feature to send to the decoder. Based on GraspCAVE, GraspTTA employs a test-time adaptation strategy for quality refinement. For SceneDiffuser, we concatenate the language and object features as the model condition. For DGTR, the language and object features are concatenated to send to its transformer decoder.
Figure 11: Visualization of grasps before and after quality grasp component. Our quality grasp component improves grasp quality and maintains intention consistency.
Figure 12: Visualization of our DexGYSGrasp framework with task-oriented simple input.
Figure 13: The illustration of our real world experiment settings.
Figure 14: Real world experiment pipeline.
Figure 15: The visualization of real world experiments.
A.4 Additional Experiments
A.4.1 Qualitative Experiments of Quality Grasp Component.
We provide additional qualitative results to verify the effectiveness of Quality Grasp Component. Figure 11 shows the grasps before and after the application of the Quality Grasp Component, demonstrating that QGC can prevent object penetration and maintain consistency with the original intention.
A.4.2 Qualitative Experiments of Task-oriented Guidance.
We conduct qualitative experiments to demonstrate the generalization of our DexGYSGrasp framework to task-oriented or functional grasp task. Specifically, we input task-oriented guidance (e.g., "use" or "hold") into our framework, which has been trained on DexGYSNet. As shown in Figure 12, our DexGYSGrasp framework exhibits good compatibility with these inputs. This further confirms that our approach enables more flexible and natural human-robot interactions.
A.5 Real World Experiments Details
Experimental Environment Figure 13 shows the settings of our real world experiments. The experiments are conducted on Allegro hand, a Flexiv Rizon4 arm and an Intel Realsense D415 camera. The experimental object is a 3D printed object from test set of DexGYSNet.
Experiment Pipeline Our DexGYSGrasp takes full point clouds as input following recent works in dexterous grasping dexgraspnet ; scene_diffuser ; xu2024dgtr . To make our methods more practical, we employ three off-the-shelf models in a cascade to obtain a full point cloud from scene point cloud. As shown in Figure 14, we input the object category and the RGB image into an open-set detection model liu2023groundingdino to detect the bounding box of the object. This bounding box is then used as a prompt for SAM kirillov2023SAM to obtain the segmentation of the object. Next, we crop the target depth image using the segmentation map and the depth input. Finally, we convert the partial depth image into point clouds and feed it into a point completion network yuan2018pcn to obtain the final full point clouds. Then, the full point clouds are fed into our framework to obtain the dexterous grasp pose, which is then transformed into the real coordinate system. In execution, we first move the arm to the 6-DOF pose of the dexterous hand root node, and then control the joint angles to achieve the target pose.
Power drill Hammer Trigger sprayer Game controller Pincer Frying pan Wineglass
Success 9/10 4/10 3/10 8/10 5/10 3/10 6/10
Table 5: The results of real word experiment.
Experiment Results The experiment results are presented in Table 5. For each object, we command robot with different language instruction, and each instruction is tested five times, resulting in a total of ten grasping trials per object. A grasp is deemed successful if it aligns with the intended instruction and maintains stability, preventing the object from falling. Our method demonstrates a moderate success rate, indicating its effectiveness. Further research on real-world scenarios is recommended to enhance the robustness of our approach.
A.6 Societal Impacts and Limitations
The core innovation of this paper has a significant positive impact on society. We propose a novel task: language-guided dexterous grasp generation, which can promote human-robot interaction and expedite the deployment of robots in real-world scenarios. Additionally, we introduce an innovative framework to accomplish this task. Our approach can generate high-quality grasps while ensuring consistency of intent and diversity of grasps.
However, our method still faces some challenges in real-world deployment. Due to limitations in the current development of robotic arm control and physical structures, we cannot guarantee success in every grasp execution in the real world. In future work, we will further enhance the quality of grasp generation to improve the success rate in real-world scenarios.